A. If you have 3.00g of H2, how many grams of NH3 can be produced?
B.How many grams of H2 are needed to react with 3.80g of N2?
C.How many gramsof NH3 can be produced from 11.6g of H2?

Nitrogen gas reacts with hydrogen gas to produce ammonia via the following reaction: N2(g) + 3H2(g)-2NH3(g) Part A If yn
Answer #1

A)

Molar mass of H2 = 2.016 g/mol

mass of H2 = 3 g
mol of H2 = (mass)/(molar mass)
= 3/2.016
= 1.488 mol


According to balanced equation
mol of NH3 formed = (2/3)* moles of H2
= (2/3)*1.488
= 0.9921 mol


Molar mass of NH3,
MM = 1*MM(N) + 3*MM(H)
= 1*14.01 + 3*1.008
= 17.034 g/mol


mass of NH3 = number of mol * molar mass
= 0.9921*17.03
= 16.9 g
Answer: 16.9 g

B)

Molar mass of N2 = 28.02 g/mol

mass of N2 = 3.8 g
mol of N2 = (mass)/(molar mass)
= 3.8/28.02
= 0.1356 mol

According to balanced equation
mol of H2 reacted = (3/1)* moles of N2
= (3/1)*0.1356
= 0.4069 mol


Molar mass of H2 = 2.016 g/mol


mass of H2 = number of mol * molar mass
= 0.4069*2.016
= 0.8202 g
Answer: 0.820 g

C)

Molar mass of H2 = 2.016 g/mol

mass of H2 = 11.6 g
mol of H2 = (mass)/(molar mass)
= 11.6/2.016
= 5.754 mol

According to balanced equation
mol of NH3 formed = (2/3)* moles of H2
= (2/3)*5.754
= 3.836 mol


Molar mass of NH3,
MM = 1*MM(N) + 3*MM(H)
= 1*14.01 + 3*1.008
= 17.034 g/mol


mass of NH3 = number of mol * molar mass
= 3.836*17.03
= 65.34 g
Answer: 65.3 g